# On the Design of One-step Diffusion via Shortcutting Flow Paths

### v3

@@ Appendix F Limitations and Future Works
  • Inflexibility in training with CFG. We observe that introducing CFG leads to a relative improvement of $(33.05-6.09)/33.05=81.5\%$, indicating that training with CFG is essential. However, the current approach follows Eq. 88, which inevitably introduces two additional hyperparameters, $\omega$ and $\kappa$. As shown in Table 5 and in Table 4 of the original work, the optimal values of these parameters, as well as the triggered intervals, vary significantly across architectures. This greatly complicates hyperparameter tuning, and for large models such as XL/2, results in substantial computational overhead. Therefore, we argue that alternative approaches, such as representation alignment (repa), representation entanglement (reen), or RL-guided generation (zheng2025diffusionnftonlinediffusionreinforcement), may offer promising replacements by injecting class-related semantic information into training or enabling CFG-free diffusion generation. We leave the exploration of these directions for future work.

### v5

@@ Appendix F Limitations and Future Works
  • Inflexibility in training with CFG. We observe that introducing CFG leads to a relative improvement of $(33.05-6.09)/33.05=81.5\%$, indicating that training with CFG is essential. However, the current approach follows Eq. 86, which inevitably introduces two additional hyperparameters, $\omega$ and $\kappa$. As shown in Table 5 and in Table 4 of the original work, the optimal values of these parameters, as well as the triggered intervals, vary significantly across architectures. This greatly complicates hyperparameter tuning, and for large models such as XL/2, results in substantial computational overhead. Therefore, we argue that alternative approaches, such as representation alignment (Yu et al., 2025), representation entanglement (Wu et al., 2025), or RL-guided generation (Zheng et al., 2025), may offer promising replacements by injecting class-related semantic information into training or enabling CFG-free diffusion generation. We leave the exploration of these directions for future work.